OpenStudy (anonymous):

log(2)(x – 2) + 5 = 8 – log(2)4 Please help!!

OpenStudy (ankit042):

hello! again I think we have done a similar problem! just take log terms on one side first

OpenStudy (anonymous):

What do you mean? Like, put them both on one side of the equation???

OpenStudy (ankit042):

yeah first bring log terms on the left side and constants on the right

OpenStudy (anonymous):

So it would turn out to be: log(2) ((x-2)*(4) = 3??

OpenStudy (ankit042):

yes good! now you know how to proceed?

OpenStudy (anonymous):

Multiply the parts to the log? so log(2) (4x-8) = 3?

OpenStudy (ankit042):

yeah then?

OpenStudy (ankit042):

you remember yesterday we learned log(a)b = c can be written as b = a^c

OpenStudy (anonymous):

Yeah, okay, so 4x-8 = 2^3??? Then we solve???

OpenStudy (ankit042):

Yes!! see you know this stuff good work

OpenStudy (anonymous):

Okay thank you XD

OpenStudy (ankit042):

XD haha ...no problem